University of Illinois Department of Economics

Spring 1996  Econ 476  Roger Koenker

Some Bibliography

Inevitably, one of the first questions raised by students in Econ 476 is, “Can you suggest some other references which might be helpful for background reading?” I am usually reluctant to do this because there is, inevitably, some question about guessing an appropriate level for the supplementary material. To circumvent this problem I have tried to assemble some references which span a reasonably wide range of levels. The references are subdivided into the two traditional categories: Probability and Statistics. Within these categories they are listed in order of approximate difficulty.

A useful overview of some of this material is contained in three books written explicitly by and for econometricians again listed in order of difficulty.

1. McCabe, B. and Tremayne, A. (1993), Elements of Modern Asymptotic Theory with Statistical Applications, Manchester U. Press.

2. White, H. (1984) Asymptotic Theory for Econometricians, Academic Press.

3. White, H. (1994) Estimation, Inference, and Specification Analysis, Cambridge University Press.

Probability

1. Chung, K.L. (1979), Elementary Probability Theory, Springer-Verlag.

2. Whittle, P. (1970), Probability via Expectation, Springer-Verlag.

3. Breiman, L. (1968), Probability (Siam reprint 1992).

4. Chung, K.L. (1974), A Course in Probability Theory, Academic Press.

5. Williams, D. (1991), Probability with Martingales, Cambridge.

6. Billingsley, P. (1979), Probability and Measure, Wiley.

7. Pollard, D. (1984), Convergence of Stochastic Processes, Springer-Verlag. Statistics

1. Bickel, P. and K. Doksum (1977), Mathematical Statistics, Holden-Day.

2. Cox, D.R. and D.V. Hinkley (1974), Theoretical Statistics, Chapman-Hall.

3. McCullagh, and P. Nelder, J.A. (1989), Generalized Linear Models, Chapman-Hall.

4. Serfling, R. J. (1980), Approximation Theorems of Mathematical Statistics, Wiley.

5. Lehmann, E. (1983), Theory of Point Estimation, Wiley.

6. Hajek, J. and Sidak, Z. (1967), Theory of Rank Tests, Academic Press.